# Tensor-Var: Variational Data Assimilation in Tensor Product Feature Space

### v2

@@ C.1 Lorenz 96
with periodic boundary conditions $x_{k+K}=x_{k}$. The first term models advection, and the second term represents a linear damping with magnitude $F$. In general, the dynamics becomes more turbulent/chaotic as F increases. We choose the number of variables $K=40,80$ and the external forcing $F=10$, where the system is chaotic with a Lyapunov time of approximately 0.6 time units. As an observation model for the following experiments, we randomly observe $25\%$ states (e.g. $10$ in $K=40$). Our models were trained on a dataset $\mathcal{D}$ of $N=100$ trajectories, each trajectory consist of $=5000$ time steps long, generated by integrating the system from randomly sampled initial conditions.
C.1.1 Data generation.
To generate the dataset, we use the 8th-order Runge-Kutta (Butcher, 1996) method to numerically integrate the Lorenz-96 systems with sample step $0.1$ and the integration step $\Delta t$ size is set to $0.01$. The system is integrated from randomly sampled initial conditions, and data is collected once the system reaches a stationary distribution. For an observation operator, we use subsampling which every 5th and 10th variable for $40$ and $80$-dimensional system are observed via the nonlinear mapping $5\arctan(\cdot\pi/10)+\epsilon$ with noise $\epsilon\sim\mathcal{N}(0,0.1)$ (see Figure 7 in Appendix C.1 for an example). The $\text{arctan}\colon\mathbb{R}\mapsto[-\frac{\pi}{2},\frac{\pi}{2}]$ squeezes the state variable $S_{k}$ into $[-\frac{\pi}{2},\frac{\pi}{2}]$, which is difficult for inverse estimation. We integrate the Lorenz96 system with observation interval $\Delta t=0.1$. The history length is set as $10$ such that $h_{t}=(o_{t-10},...,o_{t-1})$.
@@
C.1.2 Additional experiment results.
We provide qualitative results in Figure 8 for the Lorenz 96 system at two different dimensions: 40 (left) and 80 (right). Each subplot illustrates the normalized absolute error for various methods, including 3D-VAR, 4D-VAR, Frerix et al. (2021), and Tensor-Var, compared to the ground truth. The assimilation window length is set to 5 (indicated by the red dashed line), with forecasts extended for an additional 100 steps based on the assimilated results. Tensor-Var generally outperforms the other methods in both dimensions and maintains stable long-term forecasts, comparable to other model-based approaches. For 3D- and 4D-VAR with partially observed models, the observed states show minimized errors (indicated by the dark lines), while the errors of unobserved states remain uncontrollable, as clearly shown in Figure 8.
@@ C.2 Kuramoto-Sivashinsky
where $x\in[0,L]$ with periodic boundary conditions. We set the domain length $L=32\pi$, large enough to induce complex patterns and temporal chaos due to high-order term interactions (Cvitanović et al., 2010). The system state $u(x,t)$ was discretized into $n_{s}=128$ and $n_{s}=256$. The observation model is the same as in Lorenz-96, where $25\%$ states can be observed. In this case, our models were trained on a dataset $\mathcal{D}$ consisting of $N=100$ trajectories, each with $L=5000$ time steps and a discretization of $\Delta t=0.01$, sampled from the stationary distribution with different initial conditions.
C.2.1 Data generation.
To generate the dataset, we use the exponential time differencing Runge–Kutta method (ETDRK), which has proven effective in computing nonlinear partial differential equation (Cox & Matthews, 2002) with an integration step $\Delta t=0.001$ and sample step $0.01$. The system is integrated from randomly sampled initial conditions, and data is collected once it reaches a stationary distribution. For observations, we use subsampling, observing every 8th state in both 128- and 256-dimensional systems, we use subsampling which every 8th for both $128$ and $256$-dimensional system are observed with noise $\epsilon\sim\mathcal{N}(0,1)$, and $5\arctan(\cdot\pi/10)$ as nonlinear mapping (see Figure 7 in Appendix C.1 for an example). The history length is set to $10$ as well.
C.2.2 Additional experiment results.
@@ C.3 Global NWP
The data is sourced from the WeatherBench2 repository (Rasp et al., 2024). From this dataset, we randomly sample grid points with 15% spatial coverage. The sampled observations include additive noise equivalent to 0.01 times the standard deviation of the state variable, ensuring robustness against observational uncertainty (see Figure 2). For model training, we use ERA5 data from 1979-01-01 to 2016-01-01, separating data from post-2018 for testing. There were 51,100 consecutive system states with generated observations for training and 2,920 data for testing.
@@
$\theta_{u}^{i}$ and $\theta_{l}^{i}$ indicate upper and lower latitude bounds, respectively, for grid cell with latitude index $i$.
C.4 ASSIMILATION FROM SATELLITE OBSERVATIONS
C.4.1 Data generation.
We collected the weather satellite track data from https://celestrak.org/NORAD/elements/ for the period 1979-01-01 00:00:00 to 2020-01-01 00:00:00. Observations were matched to the high-resolution ERA5 dataset ($240\times 121$ grid) by identifying the nearest neighborhood grid points along the satellite track to generate observations. Furthermore, we used an observation frequency of up to every half hour within the 2 hours before each assimilation time and added white noise with a standard deviation of 1% of the standard deviation of the corresponding state variables.
C.4.2 Additional qualitative results
@@
C.5 Ablation study
In this section, we provide the details of the ablation studies of (1) the history length $m$, (2) the dimensions of feature dimension and comparison with standard kernel functions, and (3) the effects of the estimated error matrices. All ablation studies are conducted on Lorenz-96 systems with $n_{s}=40$ and $n_{s}=80$. We fixed the remaining hyperparameters consistent with the main experiments and varied only the parameters under investigation.
@@ C.5 Ablation study
We examined the impact of the history length $m$ on learning the inverse observation operator and its effect on state estimation accuracy. The feature dimensions $d_{s}$, $d_{o}$, and $d_{h}$ were held constant, while the history length was varied by adjusting the size of the final linear layer (see Table 4). The state feature dimensions $d_{s}$ were set to 60 and 120 for the two system dimensions.
@@
We implemented Tensor-Var with a Gaussian kernel using kernel PCA projected to first $60$ and $120$ eigen-coordinates in scikit-learn (Pedregosa et al., 2011) by aligning with the dimension of the used DFs. For the Gaussian kernel, we approach the pre-image problem by fitting a projection operator (Kwok & Tsang, 2004). The space $\mathcal{S}\subset\mathbb{R}^{n_{s}}$ together with the linear kernel $k(s_{i},s_{j})=s_{i}^{T}s_{j}$ forms an RKHS as well. Therefore, we can simply define the projection operator as a CME operator that maps from the feature space $\mathbb{H}_{S}$ to the original space $\mathcal{S}$ as $\hat{\mathcal{C}}_{\text{proj}}=\mathbf{S}(K_{S}+\lambda I)^{-1}\Phi_{S}^{T}$, where $\mathbf{S}=[s_{1},...,s_{n}]$. By applying $\hat{\mathcal{C}}_{\text{proj}}$ now to the state, we can obtain the mean estimation of the kernel mean embedding in the $\mathcal{S}$ such that $\hat{s}=\hat{\mathcal{C}}_{\text{proj}}\mu_{\mathbb{P}_{S}}=\mathbb{E}_{S}[\hat{\mathcal{C}}_{\text{proj}}\phi_{S}(S)]=\mathbb{E}_{S}[S]$. We normalized the dataset to a standard Gaussian distribution and select the length scale $\gamma=1.0$ by performing a cross-validation on the $\gamma=[0.5,0.75,1.0,1.25,1.5,2]$.
C.6 Model architecture
@@
Table 5: Model architecture for Global NWP with input dimension $(H,W,C)$ with $C$ physical variables and spatial resolution $H\times W$. The implementation of vision Transformer (ViT) block follows Zamir et al. (2022) with applications in DA followed by Nguyen & Fablet (2024).

### v3

@@ E.1 Lorenz 96
with periodic boundary conditions $x_{k+K}=x_{k}$. The first term models advection, and the second term represents a linear damping with magnitude $F$. In general, the dynamics become more turbulent/chaotic as F increases. We choose the number of variables $K=40,80$ and the external forcing $F=10$, where the system is chaotic with a Lyapunov time of approximately 0.6 time units. As an observation model for the following experiments, we randomly observe $25\%$ states (e.g. $10$ in $K=40$). Our models are trained on a dataset $\mathcal{D}$ of $N=100$ trajectories, each trajectory consisting of $=5000$ time steps, generated by integrating the system from randomly sampled initial conditions.
E.1.1 Data generation.
To generate the dataset, we use the 8th-order Runge-Kutta (Butcher, 1996) method to numerically integrate the Lorenz-96 systems with sample step $0.1$ and the integration step $\Delta t$ size is set to $0.01$. The system is integrated from randomly sampled initial conditions, and data is collected once the system reaches a stationary distribution. For an observation operator, we use subsampling which every 5th and 10th variable for $40$ and $80$-dimensional system are observed via the nonlinear mapping $5\arctan(\cdot\pi/10)+\epsilon$ with noise $\epsilon\sim\mathcal{N}(0,0.1)$ (see Figure 6 in Appendix E.1 for an example). The $\text{arctan}\colon\mathbb{R}\mapsto[-\frac{\pi}{2},\frac{\pi}{2}]$ squeezes the state variable $S_{k}$ into $[-\frac{\pi}{2},\frac{\pi}{2}]$, which is difficult for inverse estimation. We integrate the Lorenz96 system with observation interval $\Delta t=0.1$. The history length is set as $10$ such that $h_{t}=(o_{t-10},...,o_{t-1})$.
@@
E.1.2 Additional experiment results.
We provide qualitative results in Figure 7 for the Lorenz 96 system at two different dimensions: 40 (left) and 80 (right). Each subplot illustrates the normalized absolute error for various methods, including 3D-VAR, 4D-VAR, (Frerix et al., 2021), and Tensor-Var, compared to the ground truth. The assimilation window length is set to 5 (indicated by the red dashed line), with forecasts extended for an additional 100 steps based on the assimilated results. Tensor-Var generally outperforms the other methods in both dimensions and maintains stable long-term forecasts, comparable to other model-based approaches. For 3D- and 4D-VAR with partially observed models, the observed states show minimized errors (indicated by the dark lines), while the errors of unobserved states remain uncontrollable, as clearly shown in Figure 7.
@@ E.2 Kuramoto-Sivashinsky
where $x\in[0,L]$ with periodic boundary conditions. We set the domain length $L=32\pi$, large enough to induce complex patterns and temporal chaos due to high-order term interactions (Cvitanović et al., 2010). The system state $u(x,t)$ was discretized into $n_{s}=128$ and $n_{s}=256$. The observation model is the same as in Lorenz-96, where $25\%$ states can be observed. In this case, our models are trained on a dataset $\mathcal{D}$ consisting of $N=100$ trajectories, each with $L=5000$ time steps and a discretization of $\Delta t=0.01$, sampled from the stationary distribution with different initial conditions.
E.2.1 Data generation.
To generate the dataset, we use the exponential time differencing Runge–Kutta method (ETDRK), which has proven effective in computing nonlinear partial differential equation (Cox & Matthews, 2002) with an integration step $\Delta t=0.001$ and sample step $0.01$. The system is integrated from randomly sampled initial conditions, and data is collected once it reaches a stationary distribution. For observations, we use subsampling, observing every 8th state in both 128- and 256-dimensional systems, we use subsampling which every 8th for both $128$ and $256$-dimensional system are observed with noise $\epsilon\sim\mathcal{N}(0,1)$, and $5\arctan(\cdot\pi/10)$ as nonlinear mapping (see Figure 6 in Appendix E.1 for an example). The history length is set to $10$ as well.
E.2.2 Additional experiment results.
@@ E.3 Global NWP
The data is sourced from the WeatherBench2 repository (Rasp et al., 2024). From this dataset, we randomly sample grid points with 15% spatial coverage. The sampled observations include additive noise equivalent to 0.01 times the standard deviation of the state variable, ensuring robustness against observational uncertainty (see Figure 9). For model training, we use ERA5 data from 1979-01-01 to 2016-01-01, separating data from post-2018 for testing. There are 51,100 consecutive system states with generated observations for training and 2,920 data for testing.
Figure 9: Visualization of assimilation results for five variables from ERA5 data at time 2018-01-01 00:00. Each column (from left to right) displays the background state, observations, true state, and errors for Latent-3DVar, Latent-4DVar, and Tensor-Var. The reported error was a weighted absolute error in each pixel; see Appendix E.3 for more details.
@@
$\theta_{u}^{i}$ and $\theta_{l}^{i}$ indicate upper and lower latitude bounds, respectively, for the grid cell with latitude index $i$.
E.4 ASSIMILATION FROM SATELLITE OBSERVATIONS
E.4.1 Data generation.
We collect the weather satellite track data from https://celestrak.org/NORAD/elements/ for the period 1979-01-01 00:00:00 to 2020-01-01 00:00:00. Observations are matched to the high-resolution ERA5 dataset ($240\times 121$ grid) by identifying the nearest neighborhood grid points along the satellite track to generate observations. Furthermore, we use an observation frequency of up to every half hour within the 2 hours before each assimilation time and add white noise with a standard deviation of 1% of the standard deviation of the corresponding state variables.
E.4.2 Anomaly Correlation Coefficient Evaluation
In addition to the NRMSE results reported in Subsection 4.3, we also evaluate the assimilated fields using the Anomaly Correlation Coefficient (ACC), which measures the pattern similarity between anomalies in the analysis (or forecast) and a reference field relative to a climatological mean. This metric is widely used in numerical weather prediction to assess the accuracy of large-scale spatial patterns. We compare the ACC performance of our method with the SOTA data-driven baseline, FengWu 4D-Var (Xiao et al., 2024). ACC is formally defined as:
$\text{ACC}=\frac{\sum_{t}(\hat{s}_{t}-\bar{s})(s_{t}-\bar{s})}{\sqrt{\sum_{t}(\hat{s}_{t}-\bar{s})^{2}}\sqrt{\sum_{t}(s_{t}-\bar{s})^{2}}},$
where $\hat{s}$ is the result of assimilation, $s$ is the corresponding reference value (e.g. ERA5), and $\bar{s}$ is the climatological mean. Figure 11 reports the mean and standard deviation of ACC for five atmospheric variables over a 7-day forecast horizon, with an assimilation window of length 5. The results indicate that Tensor-Var consistently achieves higher ACC values than FengWu 4D-Var over a long horizon.
Figure 11: Comparison of ACC over a 7-day horizon for five atmospheric variables from Fengwu 4D-Var and Tensor-Var. Each time-step represents a 6-hour interval.
E.4.3 Long Roll-out Stability Test
To assess the long-term stability of our method, we conduct a one-year roll-out experiment under the following settings: spatial resolution of $240\times 121$; meteorological variables {z500, t850, q700, u850, v850}; a 6-hour time step; assimilation window length of 5 steps; and a forecast lead time of 7 days. The experiment covers the full year from January 1 to December 31, 2018, with training and evaluation settings consistent with Section 4.3.
Figure 12 illustrates the long-term behavior of Tensor-Var in comparison to FengWu 4D-Var. The results show that Tensor-Var effectively controls the estimation error over extended periods and achieves lower error magnitudes than FengWu 4D-Var throughout most of the year. However, we observe that Tensor-Var exhibits instability after approximately 800 assimilation steps, whereas FengWu maintains stable performance.
The observed instability can be attributed to the linear dynamical structure of Tensor-Var, which is able to handle short-term forecasting but can be insufficient for capturing nonlinear long-term dynamics. In contrast, the transformer-based architecture of FengWu 4D-Var appears to be more robust for extended temporal integration. These findings highlight a trade-off between assimilation efficiency and long-horizon stability. Future work will focus on improving Tensor-Var’s dynamical expressiveness while preserving its computational advantages.
Figure 12: Long-term roll-out stability test for the five meteorological variables over a one-year horizon. The plot compares the NRMSE of forecasts (dashed) and forecast-DA (solid) outputs from Tensor-Var and FengWu 4D-Var.
E.4.4 Additional Qualitative Results
@@
E.5 Ablation study
In this section, we provide the details of the ablation studies of (1) the history length $m$, (2) the dimensions of the feature dimension and comparison with standard kernel functions, and (3) the effects of the estimated error matrices. All ablation studies are conducted on Lorenz-96 systems with $n_{s}=40$ and $n_{s}=80$. We fix the remaining hyperparameters consistent with the main experiments and vary only the parameters under investigation.
@@ E.5 Ablation study
We examine the impact of the history length $m$ on learning the inverse observation operator and its effect on state estimation accuracy. The feature dimensions $d_{s}$, $d_{o}$, and $d_{h}$ remain constant, while the history length is varied by adjusting the size of the final linear layer (see Table 5). The state feature dimensions $d_{s}$ are set to 60 and 120 for the two system dimensions.
@@
We implement Tensor-Var with a Gaussian kernel using kernel PCA projected to the first $60$ and $120$ eigen-coordinates in scikit-learn (Pedregosa et al., 2011) by aligning with the dimension of the used DFs. For the Gaussian kernel, we approach the pre-image problem by fitting a projection operator (Kwok & Tsang, 2004). The space $\mathcal{S}\subset\mathbb{R}^{n_{s}}$ together with the linear kernel $k(s_{i},s_{j})=s_{i}^{T}s_{j}$ also forms an RKHS. Therefore, we can define the projection operator as a CME operator that maps from the feature space $\mathbb{H}_{S}$ to the original space $\mathcal{S}$ as $\hat{\mathcal{C}}_{\text{proj}}=\mathbf{S}(K_{S}+\lambda I)^{-1}\Phi_{S}^{T}$, where $\mathbf{S}=[s_{1},...,s_{n}]$. By applying $\hat{\mathcal{C}}_{\text{proj}}$ now to the state, we can obtain the mean estimation of the kernel mean embedding in $\mathcal{S}$ such that $\hat{s}=\hat{\mathcal{C}}_{\text{proj}}\mu_{\mathbb{P}_{S}}=\mathbb{E}_{S}[\hat{\mathcal{C}}_{\text{proj}}\phi_{S}(S)]=\mathbb{E}_{S}[S]$. We normalize the dataset to a standard Gaussian distribution and select the length scale $\gamma=1.0$ by performing a cross-validation on $\gamma=[0.5,0.75,1.0,1.25,1.5,2]$.
E.6 Model architecture
@@
Table 6: Model architecture for Global NWP with input dimension $(H,W,C)$ with $C$ physical variables and spatial resolution $H\times W$. The implementation of vision Transformer (ViT) block follows (Zamir et al., 2022) with applications in DA followed by (Nguyen & Fablet, 2024).